### v1

Abstract
Retrieval-Augmented Generation (RAG) models are critically undermined by citation hallucinations, a deceptive failure where a model confidently cites a source that fails to support its claim. Existing work often attributes hallucination to a simple over-reliance on the model’s parametric knowledge. We challenge this view and introduce FACTUM (Framework for Attesting Citation Trustworthiness via Underlying Mechanisms), a framework of four mechanistic scores measuring the distinct contributions of a model’s attention and FFN pathways, and the alignment between them. Our analysis reveals two consistent signatures of correct citation: a significantly stronger contribution from the model’s parametric knowledge and greater use of the attention sink for information synthesis. Crucially, we find the signature of a correct citation is not static but evolves with model scale. For example, the signature of a correct citation for the Llama-3.2-3B model is marked by higher pathway alignment, whereas for the Llama-3.1-8B model, it is characterized by lower alignment, where pathways contribute more distinct, orthogonal information. By capturing this complex, evolving signature, FACTUM outperforms state-of-the-art baselines by up to 37.5% in AUC. Our findings reframe citation hallucination as a complex, scale-dependent interplay between internal mechanisms, paving the way for more nuanced and reliable RAG systems.
1 Introduction
Figure 1: An example demonstrating how citation hallucination makes factually correct information dangerously unreliable. Both scenarios receive the same query and retrieved documents. Left (Correct Citation): The model correctly cites its source ‘[Source: 1]’, making the claim verifiable. Right (Hallucinated Citation): The model generates the same correct fact but falsely attributes it to ‘[Source: 2]’, pointing the user to conflicting information and making the claim unverifiable. This isolates citation hallucination as a critical failure of the model’s attribution mechanism.
Retrieval-Augmented Generation (RAG) systems are designed to improve the accuracy of Large Language Models (LLMs) by integrating relevant information retrieved from external knowledge sources [21, 2]. However, even with access to accurate information, these models can still produce statements that misrepresent or contradict the provided sources, a phenomenon known as RAG Hallucination [51, 47]. Prior work has studied the use of external context and an LLM’s parametric knowledge to detect general RAG hallucinations [47, 48, 25]. In contrast, the critical issue of citation hallucination detection has been largely unaddressed. We argue that detecting hallucination of citations in RAG is a distinct problem that deserves its own study.
Citation hallucination is the misattribution of information to an incorrect or fabricated source [28]. Figure 1 demonstrates how a hallucinated citation makes even factually correct information dangerously unreliable. The hallucination of citations is particularly problematic because the presence of citations significantly increases a reader’s trust in the presented information [8, 43]. Moreover, studies show that users are influenced by the sheer number of citations, even when the cited content fails to support the claims that were made [36]. Consequently, a model that generates hallucinated citations breaks the fundamental promise of verifiability that a reference is meant to provide [8].
This becomes particularly challenging in long-form RAG, where LLMs are tasked with generating extended, multi-paragraph reports from vast document collections [^1] Modern long-context models can process context windows from 32K to 2M tokens, equivalent to hundreds or thousands of pages of text [23].. In this scenario, the model must maintain correct attribution throughout a long output, introducing the risk of attributional drift: the longer the model writes, the more likely it is to forget which statement came from which source, leading to misattribution [6]. To identify the internal signatures that characterize hallucination of citations, our work moves beyond black-box analysis to investigate the measurable mechanistic signals within the model’s internal states.
To locate these signals, we ground our approach in the transformer’s architecture. Each layer processes information through two distinct sub-modules with fundamentally different roles. The Attention mechanism serves as the sole pathway for retrieving information from the external context, processing the provided source documents [22, 7]. In contrast, the Feed-Forward Networks (FFNs) are the primary pathway for applying the model’s stored, parametric knowledge [13, 42]. This architectural division allows us to conceptualize them as two primary information streams: what the model reads and what it recalls. By decoupling their contributions to the model’s residual stream, which is the central information space where updates are written, we can directly probe the interplay between what the model reads from the external context and what it recalls from parametric memory [9, 29].
Prior mechanistic analyses of general RAG hallucination argued that hallucinations predominantly arise from an over-reliance on the model’s parametric knowledge [47]. We argue that for the task of citation, this view is incomplete. The core failure is not a simple over-reliance on parametric knowledge, but a systemic breakdown in how the model utilizes its internal mechanisms. To reliably detect citation hallucination, we must ask a series of interconnected questions about the model’s state at the moment of citation generation. Regarding what the model reads via its Attention pathway: how strongly is the citation grounded in the source documents, and how is that information synthesized? Regarding what the model recalls from its FFNs: how forcefully does it apply its own parametric knowledge? Finally, how do these two pathways interact? Are they working in concert, or are they contributing distinct, orthogonal information?
To answer these questions, our contributions are threefold: (1) we introduce FACTUM (Framework for Attesting Citation Trustworthiness via Underlying Mechanisms), a novel diagnostic framework using refined and novel mechanistic scores to measure these signals; (2) we demonstrate that by analyzing the complete signature of citation generation, FACTUM significantly outperforms state-of-the-art baselines for citation hallucination detection in long-form RAG; and (3) we provide empirical evidence challenging the conventional view, showing that strong FFN Pathway activity is constructive, not detrimental, to correct citation generation.
2 Related Work
2.0.1 Hallucination Detection in RAG.
Methods for detecting hallucinations in RAG systems largely fall into two categories: those that treat the model as a black box, and those that analyze its internal states. Black-box approaches evaluate the model’s final output without considering its internal processing [45]. This is often done by using a powerful external LLM as a judge [41, 10], fine-tuning a separate detector model [38, 44, 40, 19], or checking for consistency across multiple sampled outputs [31]. The primary limitation of these approaches is that they are purely behavioral [16, 46]. They can flag a potential hallucination, but they cannot reveal the underlying mechanistic signals within the model that give insight into the failure. In contrast, methods that analyze model internals seek measurable signals of hallucination [39]. Early approaches focused on general uncertainty, using signals such as token probabilities and entropy [52, 17, 30]. The most recent prior work is ReDeEP, which took a more mechanistic approach [47]. It was the first to decouple the model’s reliance on external context versus its internal parametric knowledge, proposing the External Context Score (ECS) and Parametric Knowledge Score (PKS). Our work adopts this decoupling framework but offers critical advancements: we introduce more direct and robust scores for each pathway, and add two novel dimensions ReDeEP does not address: a direct measure of pathway alignment and the role of the attention sink.
While foundational, ReDeEP’s scores have critical limitations for detecting citation hallucination. First, ECS is an imprecise measure of contextual grounding for this task. As it considers the entire prompt, it dilutes the crucial signal from source documents with noise from system instructions and the user query. This design also conflates two distinct attention signals: attention to the actual evidence and attention to the frequently utilized “attention sink” (beginning-of-sentence token), making it impossible to study their separate roles. Second, PKS relies on LogitLens to project hidden states into a vocabulary space, a proxy that has been shown to be unreliable [5]. Finally, ReDeEP offers no way to directly measure the geometric alignment or conflict between the model’s information pathways. FACTUM is designed to overcome these specific limitations.
2.0.2 Citation Hallucination.
While hallucination in RAG is a broad topic, citation hallucination is a specific and critical sub-problem [8]. Recent work has established its prevalence, confirming that even state-of-the-art models frequently misattribute or fabricate sources [28]. Existing studies have approached this from a black-box perspective, for instance by asking an LLM follow-up questions to test its comprehension of a source [1, 37]. However, this still fails to answer the core mechanistic question: Is there a measurable signature within the model’s internal information pathways that reliably distinguishes a correct citation from a hallucinated citation? Our work is the first to systematically search for this internal signature, connecting citation hallucination to a predictable pattern in the model’s internal state.
Figure 2: An illustration of the FACTUM framework analyzing a model’s internal state at a citation token (e.g., 1), inspired by prior work [47]. The framework derives scores from the two primary pathways. From the Attention Pathway, it computes the Contextual Alignment Score (CAS) to measure grounding in the source documents, and the Beginning-of-Sentence Attention Score (BAS) to measure information synthesis. It also measures the FFN Pathway’s Parametric Force Score (PFS) to quantify the magnitude of the FFN’s update, and the Pathway Alignment Score (PAS) to assess the geometric alignment between the two pathway updates.
3 The FACTUM Framework
The FACTUM framework leverages the internal architecture of the Transformer to provide a direct, mechanistic view into the model’s state during citation generation. Rather than treating the model as a black box, we decompose its internal processing into two distinct information pathways: the Attention pathway (the reader) and the FFN pathway (the recall). By decoupling these updates to the residual stream, we define four mechanistic scores that are designed to overcome the limitations of prior methods by providing more direct, robust, and targeted measurements of the model’s internal processing. Together, these scores allow us to test our central hypothesis that citation hallucination arises from a systemic failure in the coordination between contextual grounding, internal synthesis, parametric force, and pathway alignment. Figure 2 provides a visual representation of these scores and how they are calculated.
FACTUM is grounded in the mechanistic interpretability view of Transformers, treating hidden states as points in a structured high-dimensional residual stream $\mathbf{x}\in\mathbb{R}^{d}$ [9]. Following prior work, we model the architecture as a shared communication channel where each sub-layer performs an additive write to that state. For a citation token at position $i$, the state at layer $l$ is the sum of the previous state and the updates from the multi-head attention (MHA) and feed-forward network (FFN) sub-layers:
$\mathbf{x}_{i}^{(l)}=\mathbf{x}_{i}^{(l-1)}+\mathbf{v}_{\text{attn},i}^{(l)}+\mathbf{v}_{\text{ffn},i}^{(l)}$ (1)
This linear, additive structure allows us to decouple the model’s internal processing into two primary information pathways, measuring their individual contributions to the communication channel [15, 29].
3.1 The Attention Pathway
The Attention Pathway serves as the model’s reader and is the sole mechanism for moving information between token positions [15, 50]. In RAG, its purpose is to contextualize the current token by selectively retrieving information from the source documents $T_{C}$ [22, 7]. Mechanistically, this pathway facilitates a targeted transport of features: it uses the QK circuit to determine the attention pattern $A_{i,j}$ (where to look) and the OV circuit ($W_{OV}$) to extract specific semantic content (what to read) [9]. The resulting update $\mathbf{v}_{\text{attn}}$ aggregates this retrieved evidence into the residual stream:
$\mathbf{v}_{\text{attn},i}^{(l)}=\sum_{j\in T_{C}}A_{i,j}^{(h,l)}(\mathbf{x}_{j}^{(l-1)}W_{OV}^{(h,l)})$ (2)
We introduce two scores to quantify how this reading mechanism is utilized to ground and synthesize information retrieved from the source documents.
3.1.1 Context Alignment Score (CAS).
To overcome the signal dilution of prior methods that consider the entire prompt, CAS provides a precise measure of contextual grounding by focusing exclusively on the source document tokens ($T_{C}$). It computes an attention-weighted average of the final-layer hidden states ($\mathbf{h}^{(L)}$) of all context tokens to form a holistic context vector, reflecting the model’s understanding of the evidence. This weighting by attention ensures that the most attended-to context tokens have the greatest influence. The score is the cosine similarity between this context vector and the generated citation token’s own final-layer hidden state:
[…]
A high CAS indicates strong semantic alignment between the generated token and the attended-to evidence, serving as a direct measure of contextual grounding.
3.1.2 Beginning-of-sentence Attention Score (BAS).
Recent work has identified the beginning-of-sentence token <s> as a functional “attention sink” utilized for information synthesis, particularly in long contexts [12]. This sink manages information flow by avoiding representational collapse and over-squashing over long contexts before making a final prediction [4, 3]. Therefore, we hypothesize that this mechanism is used constructively to integrate retrieved information before generating a citation. To measure this, BAS calculates the proportion of attention each head $h$ in layer $l$ allocates to the <s> token when generating the token citation $t_{i}$ at sequence position $i$:
$\text{BAS}^{(l,h)}(t_{i})=A_{i,1}^{(l,h)}$ (4)
A higher BAS suggests greater reliance on this internal synthesis mechanism, a signal entirely missed by previous approaches. For models without such a token, it measures attention to the very first token of the input, which serves the same functional role.
3.2 The FFN Pathway
The FFN Pathway represents the model’s recall mechanism and serves as the primary channel for injecting parametric knowledge into the residual stream [13, 7]. FFNs function as key-value memories where factual associations are stored and retrieved [11]. Following the key-value recall model, the FFN update vector at layer $l$ is defined as:
$\mathbf{v}_{\text{ffn},i}^{(l)}=\sigma(\mathbf{x}_{i}^{(l)}W_{in}^{(l)})W_{out}^{(l)}$ (5)
where $W_{in}$ acts as keys that detect semantic patterns in the communication channel and $W_{out}$ acts as values that inject the associated parametric information [11]. Crucially, direct interventions on FFNs have been shown to manipulate the factual content of a model’s output [34, 35, 24]. By measuring the activity of these memories, we can quantify the intensity of the model’s internal updates during citation generation.
3.2.1 Parametric Force Score.
To provide a more direct measure of parametric influence than unreliable proxies such as LogitLens [5], we measure the FFN’s effect on the residual stream. Since the FFN pathway applies the model’s stored factual knowledge [11, 35], their activity during citation generation can be interpreted as a form of internal fact-checking. Therefore, we hypothesize that a high PFS reflects the model confidently corroborating contextual evidence with its parametric knowledge. At each layer $l$, the FFN contributes an update to the residual stream. We isolate this contribution by taking the difference between the residual stream’s state before the FFN ($\mathbf{x}_{\text{pre-ffn}}^{(l)}$) and immediately after the FFN block ($\mathbf{x}_{\text{post-ffn}}^{(l)}$), which we call the FFN update vector ($\mathbf{v}_{\text{ffn}}^{(l)}$). The PFS is the L2 norm of this vector:
[…]
A larger norm signifies a more substantial modification to the residual stream, indicating a stronger “force” from the model’s stored knowledge.
3.3 Pathway Interaction
A key component of our hypothesis is that citation hallucination arises from a systemic coordination failure across all four of our measured signals is the direct geometric alignment between the model’s information pathways. While not the sole indicator, this alignment provides a crucial signal about whether the model’s parametric knowledge is working with or against the contextual evidence. The Pathway Alignment Score (PAS) is the first metric designed to explicitly quantify this specific interaction.
3.3.1 Pathway Alignment Score (PAS).
The PAS metric quantifies the geometric alignment between the attention and FFN pathways by computing the direct cosine similarity between their respective update vectors. This provides a rich signal that distinguishes between collaboration, conflict, and uncoordinated activity. The score is calculated as:
[…]
The attention update vector, $\mathbf{v}_{\text{attn}}^{(l)}$, is the change to the residual stream from the attention mechanism, where $\mathbf{x}_{\text{input}}^{(l)}$ and $\mathbf{x}_{\text{pre-ffn}}^{(l)}$ are the states of the residual stream before and after the attention block, respectively. The score’s value lies in its full range: scores near +1+1 indicate alignment (collaboration), scores near 0 indicate orthogonality (unrelated processing), and scores near 1-1 indicate anti-alignment where the FFN actively negates the contextual information.
4 Experimental Setup
4.1 Task Definition: Single-Citation Long-Form RAG
We focus on a common and critical RAG use case: generating a long-form report where statements are grounded in retrieved sources. More formally, given an input query $q$ and a set of retrieved source documents $D=\{d_{1},d_{2},\ldots,d_{m}\}$, the model is tasked with generating a response composed of sentences with single inline citation $c_{j}$ pointing to a document $d_{j}\in D$. The generated text follows the format: “The standard dose for adults is 100 mg daily [Source: 1].”. We define a citation hallucination as any citation not attested by its referenced document. To achieve maximum precision, we formulate this as a token-level binary classification task on the citation digit (e.g., the digit ‘1’), as this token forms the critical link to the source. While focused, this token-level approach is flexible and can be aggregated to evaluate multi-source citations.
4.2 Dataset
Our work focuses on detecting citation hallucination in a challenging long-form, multi-document RAG scenario. To this end, we use the TREC NeuCLIR 2024 dataset [^2] Dataset available at: https://neuclir.github.io/ [20]. Its report generation task, requiring synthesizing information from retrieved documents, is significantly more challenging than the single-passage QA tasks often used in prior RAG hallucination studies. In our experiments on NeuCLIR, we use 15 documents as input to the generator. This multi-document scenario creates a high risk of the “attributional drift” that our work aims to detect. Although NeuCLIR is a cross-lingual track, our analysis is performed in a monolingual English setting: we provide the models with machine-translated English versions of the source documents to generate English reports.
Crucially, while our experimental setup uses this 15-document implementation, the FACTUM framework is agnostic to the retrieval system. It operates at the moment of generation, making its scores applicable to any RAG system, regardless of the number or quality of retrieved documents.
4.3 Ground-Truth Labeling
Because our mechanistic approach requires analyzing the model’s unique generation trace, we must label the model’s own output. Given this constraint and the need for a scalable solution, we implemented an automated LLM-as-a-judge approach. Recognizing that the validity of LLM-based labels is crucial, we implemented a rigorous, multi-stage validation process for our LLM-as-a-judge (Llama-3.1-70B-Instruct):
  • Constrained task: We tasked the judge with a narrow attestation judgment using the ARGUE framework [32], not an open-ended quality evaluation. This limits subjectivity and improves reliability.
  • Stability testing: We verified the reliability of the judgments by testing for stability across multiple runs at varying temperatures. The labels were highly consistent, with less than 1.5% of citations showing any variation.
  • Human validation study: To verify the accuracy of the LLM-as-a-judge, we conducted a formal validation study using Cohen’s Kappa ($\kappa$) on a subset of 100 randomly sampled judgments (50 from Llama-3.1-8B and 50 from Llama-3.2-3B) using two independent human assessors [33]. As shown in Table 1, the human annotators achieved substantial agreement ($\kappa=0.68$). Crucially, the LLM judge demonstrated high alignment with both human raters across both model scales, confirming its reliability as a proxy for human judgment in this study.
Table 1: Validation study results using Cohen’s Kappa ($\kappa$). Scores indicate inter-rater reliability (Human-Human) and LLM-as-a-judge alignment against human ground truth.
Model Inter-Rater Human-LLM Alignment Overall
(Target) (H1 vs. H2) (LLM vs. H1) (LLM vs. H2) LLM Mean
Llama-3.1-8B 0.79 0.67 0.61 0.64
Llama-3.2-3B 0.62 0.62 0.79 0.71
Average 0.71 0.65 0.70 0.68
4.4 Models and Data Statistics
We test our framework on two state-of-the-art long-context models: Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct. The 8B model generated longer reports on average ($\approx$ 1,091 tokens) than the 3B model ($\approx$ 782 tokens). For the 3B model, our final dataset contains 208 correct and 93 hallucinated citation tokens. For the 8B model, we identified 465 correct and 107 hallucinated citation tokens. To train our classifiers, we use a balanced training set and evaluate on a held-out, imbalanced test set that mirrors the real-world distribution of correct vs. hallucinated citations.
4.5 Feature Engineering and Classification
The raw mechanistic scores from our framework produce a high-dimensional feature vector for each citation token. For the 32-layer, 32-head Llama-3.1-8B model, a single score such as CAS generates 1,024 distinct measurements (one for each head at each layer) for every citation token we analyze. The high dimensionality of these raw features relative to the number of labeled samples creates a significant risk of overfitting. To mitigate this, we developed a multi-stage pipeline to distill these raw signals into a concise and robust feature set.
First, we implement model component pruning. We hypothesize that many layers and heads are not specialized for citation attribution. Therefore, we rank all model components (i.e., each head at each layer) for each mechanistic score based on their individual correlation with the citation hallucination label. We treat the pruning percentage, $k$, as a hyperparameter to select only the top-$k\%$ model components from each of our mechanistic scores independently. This approach is agnostic to a model component’s architectural location, allowing us to discover whether the most predictive signals are distributed broadly or are localized to a specific subset of model components. Our optimization revealed that the optimal $k$ varies by model size; we report final results using the optimal $k$ for each model ($k=100\%$ for 3B, $k=25\%$ for 8B).
Second, the pruned features undergo dimensionality reduction. For attention-based scores (CAS, BAS, and the baseline ECS), we aggregate across the head dimension at each layer using simple statistics (e.g., mean, standard deviation), as heads within a layer function as an unordered set. This results in a per-layer feature vector for each score. To capture trends across the model’s hierarchical layers, we engineer features using both simple aggregates (mean, std) and sequence-aware functions (linear slope, FFT) common in time-series analysis. From this candidate pool, we select the most predictive feature for each mechanistic score using in-distribution training data. This pipeline, applied identically to our FACTUM scores and the ReDeEP baselines, yields a concise final feature set; for example, our complete FACTUM model uses just four features.
These features are used to train a suite of interpretable classifiers, including Logistic Regression [49], LightGBM [18], and an Explainable Boosting Machine (EBM) [27]. We employ 10-fold stratified cross-validation for evaluation. To prevent data leakage and ensure generalizability, all data splits are performed at the report level, guaranteeing that all citations from a given report belong exclusively to either the training or the testing set within any given fold.
4.6 Baselines
To evaluate FACTUM’s effectiveness, we benchmark its performance against state-of-the-art baselines that also rely exclusively on internal model signals [^3] Code available at https://github.com/Maximeswd/citation-hallucination. Our primary comparison is a re-implementation of the mechanistic scores from ReDeEP: the External Context Score (ECS) and the Parametric Knowledge Score (PKS) [47]. This allows for a direct, head-to-head comparison of feature effectiveness. We also compare against a range of established confidence-based scores that probe model internals for uncertainty signals. This category includes token-level Perplexity [14], LN-Entropy [30], the Energy score [26], and the P(True) score [17].
5 Results and Analysis
In this section, we first establish the superior performance of the FACTUM framework as a practical, state-of-the-art citation hallucination detector. We then leverage FACTUM’s diagnostic scores as mechanistic probes to uncover the underlying signatures of correct and hallucinated citations.
5.1 FACTUM Establishes a New State-of-the-Art in Citation Hallucination Detection
As shown in Table 2, our results demonstrate that FACTUM significantly improves feature-based citation hallucination detection. While some classifier-free baselines achieve competitive AUC scores or high recall, this performance comes at a cost to precision, making them unreliable in practice. Consider the P(True) score for instance, which achieves high recall on the 8B model. While it successfully identifies a large proportion of hallucinated citations (low false negatives), it does so at the cost of fairly poor precision. This means it incorrectly flags a vast number of correct citations as hallucinations (a high false positive rate). For any practical application, such a system would be unusable, as it would constantly undermine user trust. This highlights the critical need for a more balanced and nuanced classification approach that can reliably distinguish between correct and hallucinated citations without overwhelming the user with false alarms.
[…]
Table 2: Mean performance on the NeuCLIR 2024 dataset. The table is split into two sections: classifier-free General Baselines and our main classifier-based results. In the top section, the best general baseline for each metric is marked with a gray background. In the bottom section, highlighting is scoped within each classifier block, with the best (bold) and runner-up (underlined) performance shown. A star () indicates a statistically significant improvement over ‘ECS+PKS (Baseline)’ ($p_{FDR}<0.05$). A dagger () indicates that the improvement of ‘FACTUM (Ours)’ over ‘CAS+V_ffn (Ours)’ is statistically significant.
Classifier Method Llama-3.2-3B-Instruct Llama-3.1-8B-Instruct
AUC PCC Precision Recall F1 AUC PCC Precision Recall F1
General Baselines    Perplexity 0.638 0.085 0.373 0.739 0.489 0.670 0.137 0.237 0.629 0.334
   LN-Entropy 0.637 0.189 0.376 0.717 0.487 0.671 0.219 0.237 0.625 0.333
   Energy 0.606 0.151 0.373 0.560 0.437 0.617 0.140 0.219 0.419 0.269
   P(True) 0.481 -0.035 0.278 0.588 0.369 0.481 -0.019 0.155 0.915 0.261
EBM    ECS+PKS (Baseline) 0.610 0.180 0.391 0.583 0.463 0.613 0.151 0.242 0.571 0.337
   CAS+V_ffn (Ours) 0.668 0.272 0.429 0.587 0.488 0.687 0.255 0.278 0.660 0.388
   FACTUM (Ours) 0.700⋆† 0.341⋆† 0.473⋆† 0.591 0.515⋆† 0.736⋆† 0.326⋆† 0.322⋆† 0.669 0.432⋆†
LightGBM    ECS+PKS (Baseline) 0.588 0.149 0.388 0.553 0.450 0.549 0.070 0.212 0.534 0.302
   CAS+V_ffn (Ours) 0.645 0.237 0.426 0.604 0.491 0.592 0.124 0.236 0.583 0.333
   FACTUM (Ours) 0.682⋆† 0.302⋆† 0.436 0.624 0.507 0.693⋆† 0.269⋆† 0.300⋆† 0.652⋆† 0.408⋆†
Log. Reg.    ECS+PKS (Baseline) 0.668 0.272 0.406 0.636 0.491 0.536 0.049 0.201 0.529 0.290
   CAS+V_ffn (Ours) 0.676 0.282 0.431 0.667 0.519 0.677 0.243 0.275 0.637 0.381
   FACTUM (Ours) 0.715⋆† 0.354⋆† 0.484⋆† 0.693⋆† 0.565⋆† 0.737⋆† 0.330⋆† 0.334⋆† 0.688⋆† 0.446⋆†
5.2 Uncovering the Mechanistic Signatures of Citation Hallucination
Beyond detection, FACTUM serves as a diagnostic tool to uncover the internal signatures of citation. We initially hypothesized that correct citations would be characterized by higher contextual alignment, more utilization of the beginning-of-sentence token as an attention sink, higher parametric force, and higher pathway alignment. The results in Table 3 demonstrate that the 3B model using 100% of its model components strongly supports this: Correct citations are indeed marked by significantly higher CAS, BAS, PAS, and PFS. This indicates that at this scale, utilizing all model components, the most effective strategy for correct citation generation is one of direct reinforcement, where all internal mechanisms work in concert.
In contrast, the 8B model’s signature reveals a shift toward mechanistic specialization. Its peak performance requires pruning to its top 25% of model components, suggesting that the task is localized to a more specialized set of internal mechanisms. Within this subset, while the constructive roles of high PFS and BAS remain consistent, CAS becomes inconsistent, and PAS becomes significantly lower for our best-performing classifier (Logistic Regression). This suggests that within this specialized group of model components, the pathways are not required to simply agree, but to contribute distinct, orthogonal information to the residual stream.
This forces a refinement of our hypothesis: as models scale, and don’t use all their model components, their strategy for correct citation appears to evolve from simple “agreement” to “specialized contribution”, where pathways provide complementary, orthogonal information. FACTUM’s ability to reveal these emergent, scale-dependent strategies proves that a “one-size-fits-all” theory of citation hallucination is insufficient.
Table 3: Mechanistic signatures of correct vs. hallucinated citations. We performed a one-sided Mann-Whitney U test on the features from the ‘Full FACTUM (Ours)’ set. Cells are colored to show the direction of the significant relationship: $\uparrow$ indicates the score was significantly higher for correct citations, while $\downarrow$ indicates it was significantly lower ($p<0.05$).
Mechanistic Score Llama-3.2-3B_Instruct Llama-3.1-8B_Instruct
EBM LightGBM Log. Reg. EBM LightGBM Log. Reg.
BAS (Attention Sink Usage) $\uparrow***$ $\uparrow*$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$
CAS (Contextual Alignment) $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\downarrow***$
PAS (Pathway Alignment) $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow**$ $\uparrow**$ $\downarrow***$
PFS (Parametric Force) $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$ $\uparrow***$
*$p<0.05$, **$p<0.01$, ***$p<0.001$.
5.3 Case Study: Mechanistic Analysis of Figure 1
To illustrate the diagnostic utility of our framework, we analyze the transition from correct to hallucinated citation in Figure 1 through the interplay of our four mechanistic signals.
1) Characterizing Attributional Drift. In the correct scenario (Left), the internal states reflect a high degree of mechanistic correspondence between the claim and its origin in Document 1. In the hallucinated scenario (Right), the model’s internal signature is consistent with a vulnerability to attributional drift. Mechanistically, this can be viewed as a divergence between information streams: while the FFN pathway continues to apply significant parametric force (PFS) towards the correct factual claim (100 mg), the Attention pathway fails to ground the citation digit in the source documents (Low CAS). Furthermore, a decrease in BAS suggests a failure in the internal synthesis mechanism required to manage information flow over longer contexts. This shift indicates that the generated citation index is likely driven by parametric likelihood or local sequence patterns rather than verified contextual grounding.
2) Coordination vs. Dissonance. FACTUM identifies these failures by capturing the holistic relationship between pathways at the moment of citation:
  • Pathway Coordination (Correct Case): When the model cites Document 1, FACTUM identifies coordination across all signals. The Attention pathway retrieves the claim from the correct context (High CAS) and utilizes the attention sink for stable synthesis (High BAS). Simultaneously, the FFN pathway provides a corroborating update (High PFS) that is semantically coordinated with the retrieved evidence (High PAS). Here, the reader and recall mechanisms work in concert to support the generated token.
  • Mechanistic Dissonance (Hallucinated Case): When the model falsely cites Document 2, a systemic dissonance emerges. Because Document 2 contains conflicting information (10 mg), the Attention pathway exhibits low semantic alignment with that source (Low CAS) and reduced synthesis stability (Low BAS). Crucially, the FFN pathway continues to prioritize the “100 mg” fact (Sustained PFS), resulting in an update that is orthogonal or contradictory to the cited context (Low/Divergent PAS).
By identifying this internal dissonance, FACTUM serves as a potential verifiability guardrail. As Table 3 demonstrates, correct citations are rarely defined by a single metric; rather, it is the systemic coordination of grounding, synthesis, force, and alignment that differentiates verifiable attribution from mechanistic drift, a signature entirely invisible to black-box evaluation methods.
6 Conclusion
We introduced FACTUM, a framework that establishes a new state-of-the-art method for detecting citation hallucination by analyzing a model’s internal pathways. As detailed above, its key contribution is revealing the evolving, scale-dependent signatures of correct attribution. These findings challenge prevailing assumptions about the role of parametric knowledge and demonstrate the need for nuanced, mechanistic tools to build more reliable and trustworthy RAG systems.
A Final Note on Trust of Citations.
A question for the reader: Did you notice our citation of Shakespeare in the introduction to support a claim about how citations incite trust? Although the Sonnets explore themes of trust, making the reference seem plausible, it was a deliberate false citation. Now imagine if the citation appeared as: “…incite trust in the reader [43]*”. That asterisk would represent a potential fix: a visual warning generated by a system like FACTUM, indicating the citation’s grounding is highly suspect.
We included this small deception to make the central point of this paper tangible: even for experts reading a paper on this exact topic, the pull to trust a citation is powerful. This subtle gap between plausibility and verifiable truth is precisely highlighting the critical need for our robust, mechanistic detection method FACTUM.
Acknowledgments
This research was supported by project VI.Vidi.223.166 of the NWO Talent Programme, which is (partly) financed by the Dutch Research Council (NWO).